2. A producer of pens has fixed costs of $36,000 per month which are allocated to the operation and variable costs are $1.60 per pen.

(a) Find the break-even quantity if pens sell at $2.0 each.

(b) Find the profit/loss if the company produces 65,000 pens and pens sell at $2.0 each?
Answer #1

Answer

Given information:

Fixed cost= $36000

Variable cost= $1.60/ pen

Revenue per pen= $2

a) calculation of break even quantity: = fixed cost/(RV-VC)

= 36000/(2-1.6)

= 90000

b) if the company produces 65000 pens and sells it at $2 each,

Then, there would be a loss, because, break even quantity is 90000.

Therefore,

90000*2=180000

65000*2= 130000

Difference is =( 50000) loss